OpenStudy (anonymous):

Two angles are complementary. The sum of the measure of the first angle and one-fourth the second angle is 73.5 degrees. Find the measure of the angles. Measure of the smaller angle is ?? and the measure of the other angle is??

OpenStudy (anonymous):

that is exactly how the question is worded. I double-checked.

OpenStudy (anonymous):

Thanks anyway...I'll skip this one and try it later.

OpenStudy (anonymous):

x+(1/4)y=73.5 x+y=90 now solve

OpenStudy (anonymous):

If this involves multiplying a fraction, I am lost already. Fractions are my enemy.

hero (hero):

Oh, yeah, that makes sense

OpenStudy (anonymous):

LOL...to YOU maybe

hero (hero):

Well, its just two sentences that use the same variable. The first sentence says that the two angles are complementary, meaning they add to 90 degrees. You use x and y to represent both angles and their sum adding to 90 degrees. The second sentence says take a fourth of whatever the the second angle is, then add it to the first angle, and you will get 73.5 degrees.

OpenStudy (anonymous):

Yea, but there are many different numbers you could use to make x + y = 90. How do you know which ones to use?

OpenStudy (anonymous):

73.5/4 would be 18.375 as you said earlier.

hero (hero):

No, you can't solve it like that. Forget what I did.

OpenStudy (anonymous):

okay

hero (hero):

This is what is called a systems of equations problem. You have to solve both equations for y, then set the equations equal to each other, then solve for x: y = 90 - x y = 4(73.5 - x) 90 - x = 4(73.5 - x) 90 - x = 294 - 4x 3x = 204 x = 68 y = 22

OpenStudy (anonymous):

I am trying to understand what you did...

OpenStudy (anonymous):

But I see how 68 and 22 = 90

OpenStudy (anonymous):

I need to study what you did a little longer. So I am assuming the smaller angle is 22 degrees and the larger is 68 degrees?

hero (hero):

1. The original equations were: x + y = 90 (because the angles are complementary) x + y/4 = 73.5 (because the second angle is only 1/4 of y and adds with x to get 73.5) 2. Solve both equations for y: y = 90 - x y = 4(73.5 - x) 3. Set y = y ( when you do the this, the first equation is set equal to the second and also the variable y is eliminated): 90 - x = 4(73.5 - x) 4. From here you just solve for x:

hero (hero):

Had to fix something...

OpenStudy (anonymous):

I think I am getting it slowly but surely. Thanks very much.

hero (hero):

I can post a classroom to explain it if you want...

OpenStudy (anonymous):

If I had the time I certainly would. I only have 3 hours to get this turned in and I only have 6 questions answered...

hero (hero):

I can also help you with the other questions

OpenStudy (anonymous):

okay, well that sounds good. Lord knows I can use the help. LOL...where do I go?

hero (hero):

What is your email address? I can send you a link to the classroom. Once I have your email you can delete it.

hero (hero):

This is so that pranksters don't try to join the class as well.

hero (hero):

Okay, you can delete it now.

OpenStudy (anonymous):

Did you send the link? It's not there yet.

hero (hero):

Sorry, it takes a minute to set up the classroom. I hope you have a microphone or headphones

OpenStudy (anonymous):

I don't know if I do or not. I don't have headphones and have never used the mic

hero (hero):

Hmmm, well, we won't be able to do this then. You don't have speakers either?

OpenStudy (anonymous):

Yes, I have speakers

hero (hero):

Okay, well, you'll have to use those then. Maybe you can also get your mic to work

OpenStudy (anonymous):

I may have to do this later anyway...my son is hungry and wants me to fix him something to eat.

OpenStudy (anonymous):

I will try to get back soon. I have to get this turned in tonight!!

hero (hero):

I will post the link either way when I'm done setting it up

hero (hero):

Check your email

OpenStudy (anonymous):

ok, thanks

hero (hero):

Okay, you'll have to post your email again, sorry. I had to completely re-set up my account and I didn't save your email sorry.

hero (hero):

I have the link now.